OpenStudy (anonymous):

Need Help in midterm q4, and q1 (I1,I2 Power ,last one)

OpenStudy (anonymous):

values??

OpenStudy (anonymous):

q4 VS=100V, R=5kΩ, and the transistor parameter f=1mAV3. The amplifier is biased with VIN=2V

OpenStudy (anonymous):

@hrene thatz Q5

OpenStudy (nurali):

question 4 1. (R*b)/(1+R*b)*vIN 2. (VS*(1+R*b)-vIN)/(1+R*b) question 5 Vgs=Vin 1) Vout=Vs-(f*(Vgs^3))*R 2) Gain = -3*f*(vgs^2)*R

OpenStudy (anonymous):

ohh sorry,

OpenStudy (anonymous):

vout = 800

OpenStudy (anonymous):

thnxx nurali and ali

OpenStudy (anonymous):

sorryyyy wrong answer